OpenStudy (anonymous):

State the probability of the given event for one roll of a number cube with faces labeled 1– 6. Write the probability in words, as a fraction, as a decimal, and as a percent. 1.an even number 2.a 6 3.a number less than 4 4.a number greater than 6 5.a 2 or 5 6.an odd number

Directrix (directrix):

There are 6 outcomes when a single die is tossed. 1, 2, 3, 4, 5, or 6

Directrix (directrix):

I'll help you get started with #1, 1.an even number Of these 1, 2, 3, 4, 5, or 6, the even numbers are 2, 4, and 6

Directrix (directrix):

Probability is the number of ways to get what you want to happen divided by the number of ways that, in this case, the die could fall. P(2,4,6) = 3/6 =1/2 As a decimal, 1/2 = .5 As a percentage, .5 = 50%.

OpenStudy (anonymous):

2) a 6 If there's 6 numbers on a number cube, how many 6s are there?

OpenStudy (anonymous):

You're probability formula is 

OpenStudy (anonymous):

So you only want the number 6, which is the desired outcomes and you know there's only one six on the number cube and you also know there's 6 POSSIBLE different outcomes. So can you figure out the probability is as a fraction from that?